# Memory-Efficient Personalization of Text-to-Image Diffusion Models via Selective Optimization Strategies

### v1

@@ Abstract
These authors contributed equally.$\dagger$ Work done during an internship at Qualcomm AI Research.$\ddagger$ Qualcomm AI Research is an initiative of Qualcomm Technologies, Inc.
@@ 1 Introduction
Figure 1: Key observations for selective optimization. (a) MeZO requires target-specific knowledge; applying it from scratch fails to personalize the model, while using it after partial fine-tuning improves personalization. (b) Fine-tuning results for BP-high and BP-low trained separately at different timestep ranges. BP-high captures and retains personalization features across all timesteps, whereas BP-low suffers from structural distortions due to resolution mismatch, especially in higher timestep ranges (750–1000).
@@ 2.3 MeZO on High-Resolution Images
We use MeZO with gradient accumulation to improve gradient approximation while maintaining memory efficiency. For each original image $x$ (e.g., $512\times 512$), MeZO computes the accumulated gradient. For each perturbation $n\in\{1,\dots,N\}$, we sample a random perturbation vector $\boldsymbol{z}^{(n)}\sim\mathcal{N}(0,I_{d})$ from a standard normal distribution and compute the gradient estimate:
@@
Figure 2: Comparison of TAP and DTAP over training steps when $t_{\text{max}}=1000$, $i_{\text{max}}=1000$, $k=0.05$, and $t_{\text{mid}}=750$. The black dashed line represents the Timestep-Aware Probability (TAP), where $t_{\text{mid}}$ remains fixed throughout training. In contrast, the colored curves illustrate the Dynamic Timestep-Aware Probability (DTAP) at different training steps $i$.
@@ 2.4.2 Dynamic Timestep-Aware Probability (DTAP)
where $t_{\text{start}}=t_{\text{max}}$ and $t_{\text{end}}=2t_{\text{mid}}-t_{\text{max}}$. At $i=0.5i_{\text{max}}$, $t_{\text{dyn}}(i)=t_{\text{mid}}$, aligning DTAP with TAP.
The final probability becomes:
$p^{zo}_{i,t}=\frac{1}{1+\exp^{-k(t-t_{\text{dyn}}(i))}}.$ (5)
@@
1:Pretrained diffusion model $\theta_{0}$, personalization images $\mathcal{X}$, prompt $c$, total number of training steps $i_{\text{max}}$, maximum timestep $t_{\text{max}}$, number of perturbations $N$
@@
6:     if random() $<p^{zo}_{i,t}$ then $\triangleright$ BP-low
@@ 4 Conclusion
We presented a memory-efficient optimization framework for on-device personalization of text-to-image diffusion models. Rather than naively combining low-resolution backpropagation and zeroth-order optimization, our method adaptively selects between them based on key observations of diffusion timesteps. This improves personalization and structural consistency, and highlights the potential of resolution-aware training and gradient-free optimization in diffusion models. We hope this encourages further exploration of efficient personalization strategies.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Figure 1: Key observations on different optimization strategies. (a) ZO-high requires target-specific knowledge; applying it from scratch fails to personalize the model, while using it after partial fine-tuning improves personalization. (b) Comparison of fine-tuning results across resolutions and timestep ranges using backpropagation. BP-low suffers from structural distortions due to resolution mismatch between training and inference, especially in higher timestep ranges, while BP-high achieves strong personalization but is memory-intensive and impractical for edge devices.
@@ 2.3 MeZO on High-Resolution Images
We use MeZO with gradient accumulation to improve gradient approximation while maintaining memory efficiency. For each original image $x$ (e.g., $512\times 512$), MeZO computes the accumulated gradient. For each perturbation $n\in\{1,2,\dots,N\}$, we sample a random perturbation vector $\boldsymbol{z}^{(n)}\sim\mathcal{N}(0,I_{d})$ from a standard normal distribution and compute the gradient estimate:
@@
Figure 2: Comparison of TAP and DTAP over training steps when $t_{\text{max}}=1000$, $i_{\text{max}}=1000$, $k=0.05$, and $t_{\text{mid}}=750$. The black dashed line represents the Timestep-Aware Probability (TAP), where $t_{\text{mid}}$ remains fixed throughout training. In contrast, the colored curves illustrate the Dynamic Timestep-Aware Probability (DTAP) at different training steps ii.
@@ 2.4.2 Dynamic Timestep-Aware Probability (DTAP)
where $t_{\text{start}}=t_{\text{max}}$ and $t_{\text{end}}=2t_{\text{mid}}-t_{\text{max}}$. At $i=0.5i_{\text{max}}$, $t_{\text{dyn}}(i)=t_{\text{mid}}$, aligning DTAP with TAP. The final probability becomes:
$p^{zo}_{i,t}=\frac{1}{1+e^{-k(t-t_{\text{dyn}}(i))}}.$ (5)
@@
1:Pretrained diffusion model $\theta_{0}$, personalization images $\mathcal{X}$, prompt $c$, total number of training steps $i_{\text{max}}$, maximum timestep $t_{\text{max}}$, number of perturbations NN
@@
6:  if random() $p^{zo}_{i,t}$ then $\triangleright$ BP-low
@@ 4 Conclusion
We presented a memory-efficient optimization framework for on-device personalization of text-to-image diffusion models. Rather than naively combining low-resolution backpropagation and zeroth-order optimization, our method adaptively selects between them based on key observations of diffusion timesteps. This improves personalization and structural consistency, and highlights the potential of resolution-aware training and gradient-free optimization for diffusion models. We hope this encourages further exploration of efficient personalization strategies.